Grade 12Mechanics

What is the centre of mass of hemi hollow share .derive it`s formula

Vijaykumar sonu
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
Dear Student,
For small element dm,
  1. dm= M/2πR2 x 2πRsinα Rdα
    =Msinα dα

    Yc= 1/Mꭍdmy
    =1/M (from 0 to π/2)ꭍMsinαdαRcosα
    =-R/2[cos2α/2] (from 0 to π/2)
    =R/2.
    Cheers!!
    Regards,
    Vikas (B. Tech. 4th year
    Thapar University)